# When Do LLMs Admit Their Mistakes? Understanding the Role of Model Belief in Retraction

### v2

@@ Abstract
Can large language models (LLMs) admit their mistakes when they should know better? In this work, we define the behavior of acknowledging errors in previously generated answers as “retraction” and aim to understand when and why LLMs choose to retract. We first construct model-specific datasets to evaluate whether a model will retract an incorrect answer that contradicts its own parametric knowledge. While LLMs are capable of retraction, they do so only infrequently. We demonstrate that retraction is closely tied to previously identified indicators of models’ internal belief: models fail to retract wrong answers that they “believe” to be factually correct. Steering experiments further demonstrate that internal belief causally influences model retraction. In particular, when the model does not believe its answer, this not only encourages the model to attempt to verify the answer, but also alters attention behavior during self-verification. Finally, we demonstrate that simple supervised fine-tuning significantly improves retraction performance by helping the model learn more accurate internal beliefs. Code and datasets are available on https://github.com/ayyyq/llm-retraction.
@@ 1 Introduction
Figure 1: Refer to caption indicates a correct answer, Refer to caption indicates a wrong answer, and Refer to caption denotes a retraction. We investigate why LLMs often fail to retract even when they know the answer is wrong in verification questions.
Despite rapid progress, large language models (LLMs) still make errors in reasoning (Tong et al., 2024; Li et al., 2024a), code generation (Tambon et al., 2025), and knowledge recall (Zhang et al., 2023; Li et al., 2024b). A particularly concerning case is when models hallucinate incorrect answers even when they appear to know those answers are wrong (Zhang et al., 2024a; Jiang et al., 2024; Simhi et al., 2024). For such hallucinations, an ideal response would be to promptly recognize the error and indicate it is incorrect—an act we define as retraction, as illustrated in Figure 1. Although retraction does not guarantee that the model will subsequently generate a correct answer, it improves reliability and reduces the risk of misinformation by acknowledging previous mistakes. Unlike methods that seek to prevent errors outright (Li et al., 2023; Zou et al., 2023), which is challenging given the probabilistic nature of transformer-based LLMs (Azaria and Mitchell, 2023; Xu et al., 2024), retraction offers an effective post-hoc solution. In this work, we focus on knowledge-related factoid questions and aim to understand when and why LLMs autonomously retract incorrect answers that they know are incorrect.
Given the lack of a suitable testbed for studying retraction, we first construct model-dependent “continuation” datasets. We obtain two datasets of knowledge-based questions on which LLMs often hallucinate: a set of constraint satisfaction questions where each question includes two constraints that the answer must satisfy to be correct (e.g., “Name a politician who was born in New York City”, Dhuliawala et al., 2024; Yüksekgönül et al., 2024), as well as a set of “reversal curse” questions that ask for a celebrity given their lesser-known parents (Berglund et al., 2024). Since these datasets induce many LLM errors, they provide an ideal testbed for studying retraction. For each question, we collect the target model’s self-generated incorrect answers and keep only those where the model disagrees with its initial answer when separately asked a verification question. In this way, we build continuation datasets consisting of question-answer pairs, and then prompt the model to continue generating text after the incorrect answer to see whether it will retract. While models sometimes retract their own incorrect answers, they are generally reluctant to do so, even when they have the requisite knowledge.
@@
Next, we show that a model’s decision to retract is closely linked to its internal belief about the answer’s correctness. Prior work has found directions in LLM hidden states that represent models’ internal beliefs about whether a statement is factually correct (Li et al., 2023; Liu et al., 2024). Through probing experiments, we show that these internal beliefs are indicative of models’ retraction decisions: models tend to believe that non-retracted answers are “correct” and retracted ones are “incorrect,” even when this does not align with ground truth. This link is causal: steering the model to believe an answer is correct (“positive” belief steering) suppresses retraction, while steering it to believe an answer is incorrect (“negative” belief steering) strongly promotes retraction. By analyzing the steered models, we identify two separate pathways through which internal beliefs control retraction. Steering in the negative belief direction first encourages the model to try generating additional information (e.g., the birthplace of the entity) for verification, rather than stopping immediately after the answer. Then, negative belief steering increases the model’s attention to the given answer and refines the answer’s attention value vectors, which further promotes retraction.
@@
Finally, we show that straightforward supervised fine-tuning (SFT; Prakash et al., 2024) can greatly improve in-distribution retraction performance—the model retracts more incorrect answers while still committing to correct ones. Our findings connecting model belief with retraction generalize to fine-tuned models, as the original belief direction continues to influence retraction behavior after fine-tuning. SFT works by aligning the model’s internal beliefs with ground truth, leading to more accurate retraction decisions. These fine-tuning results highlight the potential of diverse, retraction-focused training data to create LLMs that robustly retract their incorrect answers.
@@
To summarize, our contributions are as follows: (1) We construct a model-specific testbed to evaluate an LLM’s retraction performance, and show that current LLMs can retract but do so only rarely. (2) We uncover a connection between a model’s internal belief and its external retraction behavior, and identify the underlying mechanism that governs this behavior. (3) We demonstrate that the causal influence of internal belief on retraction generalizes to supervised fine-tuned models, where more accurate beliefs lead to improved retraction performance.
@@ 2 Related Work
2.1 Probing LLMs’ Belief
The beliefs of LLMs refer to their internal judgments about the truth of the world (Levinstein and Herrmann, 2023; Schouten et al., 2024). Prior work typically deciphers these beliefs by probing internal representations (Azaria and Mitchell, 2023; Marks and Tegmark, 2023a; Li et al., 2023; Burns et al., 2023). Most relevant to our work, Liu et al. (2024) propose the existence of a universal truthfulness hyperplane that can separate true and false statements based on a model’s internal representations.
We note that these probing methods may not always accurately distinguish true and false statements. Many studies focus on distinguishing synthetically constructed true-false claims, which may not reflect the distribution of hallucinations in real LLM outputs. More critically, while some research demonstrates strong performance in detecting hallucinations on in-distribution, model-generated data (Azaria and Mitchell, 2023; CH-Wang et al., 2024; Orgad et al., 2024), they fail to generalize to out-of-distribution examples (Levinstein and Herrmann, 2023; Liu et al., 2024). Our work provides a possible explanation: low out-of-distribution probe accuracy may reflect the need for a finer-grained classification of incorrect answers, such as distinguishing between retracted and non-retracted ones.
@@ 2.2 Self-Correction in LLMs
Retraction marks an important step of self-correction, in which a model must first identify its mistake, then subsequently produce a correct answer or a more accurate reasoning step. While there has been debate over whether LLMs can truly self-correct in reasoning (Huang et al., 2024; Tyen et al., 2024), it has been demonstrated that self-correction is possible and exceptionally effective when the model has necessary knowledge (Dhuliawala et al., 2024). However, previous work on self-correction primarily relies on multi-turn procedures, such as asking the model verification questions (Dhuliawala et al., 2024; Wu et al., 2024), prompting it to give feedback (Madaan et al., 2023; Zhang et al., 2024b; Liu et al., 2023), or directly instructing it to verify its initial responses (Kadavath et al., 2022; Yang et al., 2024a). These approaches are not fully automatic. In contrast, our focus is to evaluate and understand LLMs’ capabilities to autonomously identify and admit their own mistakes, thus initiating the self-correction process without explicit external prompting.
@@ 3.1 Task Definition
Retraction denotes a model’s immediate acknowledgment that its generated answer is incorrect or does not fully satisfy the user’s requirements, regardless of whether it later produces the correct answer. To evaluate the retraction performance of current LLMs, we construct a model-specific testbed. We first collect questions from two knowledge-related datasets, Wikidata (e.g., “Name a writer who was born in Oran, Algeria”) and Celebrity (e.g., “Name a child of Joe Jackson”), which are prone to eliciting wrong answers, thereby creating a great opportunity to study retraction. Details of these two original datasets are provided in Appendix B.1.
Continuation Dataset Construction.
Based on the collected questions, we construct model-specific “continuation” datasets. Each example pairs a question with a model-generated answer, and we prompt the model to continue generating to evaluate whether it will retract, as illustrated below:
@@
To ensure that each incorrect answer is, in principle, correctable by the target LLM, we first collect the model’s freely generated answers using temperature sampling. For each answer, we create verification questions (e.g., Where was {model’s answer} born? What is {model’s answer}’s profession?) and assess whether the model’s responses to these questions conflict with the constraints of the original question. We retain two types of examples:
@@
  • Wrong Examples: The answer is factually incorrect, and the model’s responses to the verification questions contradict the original question.
Correct examples are used to evaluate over-retraction and for in-distribution SFT in Section 6. We also create train/test split: all results are evaluated on the test set, while the training set is used for SFT in Section 6. We conduct experiments using three popular LLMs from different model families, Llama3.1-8B-Instruct (Dubey et al., 2024, abbr. Llama3.1-8B), Qwen2.5-7B-Instruct (Yang et al., 2024b, abbr. Qwen2.5-7B), and Olmo2-1124-7B-Instruct (OLMo et al., 2025, abbr. Olmo2-7B). The data statistics are listed in Table 1. In the following sections, we use Wikidata and Celebrity to denote the model-specific continuation datasets instead of the original datasets.
Llama3.1-8B Qwen2.5-7B Olmo2-7B
# Train # Test # Train # Test # Train # Test
Wikidata 1934 1202 1496 1072 1796 1260
Celebrity 1550 826 1142 1209
Table 1: Continuation dataset statistics. Note that Qwen2.5-7B and Olmo2-7B do not have training sets for Celebrity due to an insufficient number of correct examples. See Appendix B.2 for details.
@@
We use Llama3.3-70B-Instruct [^1] https://huggingface.co/meta-llama/Llama-3.3-70B-Instruct as a judge (Zheng et al., 2023) to automatically assess whether the target model retracts the given answer in its response. See Appendix B.3 for details. We then calculate the following two metrics to evaluate the model’s retraction performance:
$\text{Retraction Recall}=\frac{|\text{Wrong \& Retraction}|}{|\text{Wrong}|},\qquad\text{Retraction Precision}=\frac{|\text{Wrong \& Retraction}|}{|\text{Retraction}|}.$
@@ 3.2 Models Can Retract, but Do So Infrequently
Llama3.1-8B Qwen2.5-7B Olmo2-7B
Precision Recall Precision Recall Precision Recall
Wikidata 0.9012 0.2529 0.8824 0.1119 0.9881 0.1317
Celebrity 0.7722 0.1477 0.9667 0.0290 0.8824 0.0150
@@
As shown in Table 2, models consistently have low but non-zero retraction recall on our datasets. We infer that LLMs have the capability to retract incorrect answers, but the consistently low recall (at most $25\%$) highlights that such retractions are rare. Recall that our verification questions provide clear evidence that the model knows that the incorrect answers in our datasets are indeed incorrect. Thus, it appears the model has both the knowledge and the ability to retract. Then, why do LLMs fail to retract more incorrect answers? What factors govern their retraction behavior?
@@ 4 Model Belief Guides Retraction
To understand why LLMs are often unwilling to retract their own incorrect answers, we first investigate whether LLMs actually know that incorrect answers are incorrect in the context of answering the question. To explore this, we identify a universal belief hyperplane and examine whether the model’s internal beliefs align with factual correctness. By belief, we refer to the model’s internal assessment of whether an answer is correct or incorrect, which may not always match the ground truth.
4.1 Probing for Internal Beliefs
@@
Following prior work (Azaria and Mitchell, 2023; Marks and Tegmark, 2023a; Li et al., 2023), we infer a model’s internal beliefs through probing. Liu et al. (2024) show that training on a diverse set of datasets can reveal a universal belief hyperplane, rather than overfitting to in-distribution patterns. Motivated by this, we use a subset of the dataset collection from Liu et al. (2024) to train our probes, including 800 examples each from Natural Questions (Kwiatkowski et al., 2019), Trivia QA (Joshi et al., 2017), and SciQ (Welbl et al., 2017). These datasets are all short-answer closed-book QA tasks and share a similar format with Wikidata and Celebrity. Each dataset includes a 50/50 split of correct and incorrect answers, with the incorrect answers generated by GPT-4-turbo. We denote this collection as Universal Truthfulness QA dataset (UTQA).
@@
Figure 2: Average probe scores across layers for Llama3.1-8B, grouped by factual correctness and retraction behavior. Higher scores indicate that the model internally believes the answer to be correct.
For each layer of the target LLM like Llama3.1-8B, we train a separate linear probe on the UTQA training set based on that layer’s hidden states after the given answer. These probes learn to distinguish correct and incorrect answers on UTQA, and thus serve as a proxy for the model’s internal beliefs: lower prediction scores indicate the model believes the answer is incorrect, while higher prediction scores suggest it believes the answer is correct. We then apply the probes to examples from Wikidata and Celebrity, group them according to the answer’s factual correctness and the model’s retraction behavior, and plot the prediction scores across layers. The four groups are defined as follows:
  • CN: the answer is factually correct and the model does not retract it.
  • CR: the answer is factually correct and the model retracts it.
  • WR: the answer is wrong (factually incorrect) and the model retracts it.
  • WN: the answer is wrong (factually incorrect) and the model does not retract it.
From Figure 2, we find that these belief probes assign high scores to CN and WN examples, and low scores to CR and WR examples. This indicates that the model’s beliefs do not align with factual correctness, but instead align more closely with its retraction behavior: low belief scores correspond to retraction, while high scores correspond to non-retraction. For example, WN examples contain factually incorrect answers, yet the probes assign them high belief scores, implying that the model believes these answers to be correct. A similar trend is observed for Qwen2.5-7B and Olmo2-7B, as shown in Appendix C.1.
4.2 Steering Internal Beliefs Affects Retraction
Our probing results establish a correlation between the model’s internal beliefs and its retraction behavior. To demonstrate that internal beliefs causally influence retraction behavior, we steer the model’s hidden states towards belief+ (i.e., believe an answer is correct) and belief- (i.e., believe an answer is incorrect) directions.

### v3

@@ Abstract
Can large language models (LLMs) admit their mistakes when they should know better? In this work, we study when and why LLMs choose to retract, i.e., spontaneously and immediately acknowledge their errors. Using model-specific testbeds, we find that while LLMs are capable of retraction, they do so only rarely, even when they can recognize their mistakes when asked in a separate interaction. We identify a reliable predictor of retraction: the model’s momentary belief, as measured by a probe on its internal states that is trained to predict correctness on external datasets unrelated to retraction. A model retracts only when it “believes” its answers to be incorrect during generation; these beliefs frequently diverge from models’ parametric knowledge as measured by factoid questions. Steering experiments further demonstrate that model belief causally drives retraction. In particular, when the model believes its answer to be incorrect, this not only encourages the model to attempt further verification, but also alters attention dynamics. Finally, we show that supervised fine-tuning improves retraction performance by helping the model learn more accurate internal belief. Code and datasets are available on https://github.com/ayyyq/llm-retraction.
When Do LLMs Admit Their Mistakes?
Understanding The Role Of Model Belief In Retraction
Yuqing Yang University of Southern California yyang063@usc.edu          Robin Jia University of Southern California robinjia@usc.edu
@@ 1 Introduction
Figure 1: Refer to caption indicates a correct answer, Refer to caption indicates a wrong answer, and Refer to caption denotes a retraction. We investigate when LLMs fail to retract, even when they know the answer is wrong in verification questions.
Despite rapid progress, hallucinations (Zhang et al., 2023; Kalai et al., 2025) remain a fundamental challenge for current large language models (LLMs), even when they appear to have relevant parametric knowledge (Zhang et al., 2024a; Jiang et al., 2024; Simhi et al., 2024). Beyond preventing such potentially correctable errors outright (Li et al., 2023; Zou et al., 2023), an alternative post-hoc remedy is when a model, after hallucinating, spontaneously recognizes and acknowledges its mistake—an act we define as retraction, as illustrated in Figure 1. When done by an LLM without external prompting, retraction reduces the user’s burden of interrogating the model, while mitigating misinformation risk and enhancing reliability. In this work, we focus on knowledge-related questions and investigate the retraction behavior of LLMs, asking when and why models autonomously retract incorrect answers that they should know are incorrect. [^1] We do not study large reasoning models (e.g., DeepSeek-AI et al., 2025), which frequently retract (e.g., “Wait, no, that’s not right…”) but also redundantly explore alternative answers (Chen et al., 2024). Instead, we focus on standard LLMs, aiming to inspire research on mitigating hallucinations through retraction when fast answers are desired.
To study this, we build model-specific testbeds called continuation datasets. In these datasets, the model first produces an answer and then is prompted to continue generating text to see whether it retracts. We rely on two datasets that are more likely to induce hallucinations: (1) Wikidata, which requires satisfying two conditions for correctness (e.g., “Name a politician who was born in New York City; Dhuliawala et al., 2024), and (2) Celebrity, which asks for a celebrity given their lesser-known parents (Berglund et al., 2024). For each question, we collect incorrect model answers and focus on the cases where the model’s responses to verification questions (e.g., “Where was Hillary Clinton born?”; Dhuliawala et al., 2024) indicate that it knows the answer is incorrect. This yields a continuation dataset of question-answer pairs. We find that models do sometimes retract their own incorrect answers, but they are generally reluctant to do so despite having the requisite knowledge.
This raises the question: Why do models fail to retract in these cases? Prior work has used probes on models’ hidden states to infer their internal beliefs about whether a given statement is factually correct (Azaria and Mitchell, 2023; Li et al., 2023; Liu et al., 2024). This leads to two hypotheses: (1) Models may internally believe that their wrong answers are true, which causes them to not retract, or (2) Models may recognize their answers to be false, yet still choose not to verbalize this belief. We find that (1) is correct. Despite being trained to predict factual correctness, internal belief probes cannot distinguish between correct and incorrect answers during generation on our datasets. Notably, this implies that models’ “momentary” beliefs during generation can contradict their parametric knowledge elicited by verification questions, providing further evidence of LLMs’ weakness in manipulating stored knowledge (Allen-Zhu and Li, 2025). On the other hand, internal belief probes are much better indicators of whether the model will retract: models tend to retract answers they internally believe are wrong and commit to those they believe are correct.
@@
We further show that this link is causal: steering the model to believe an answer is correct (positive belief steering) suppresses retraction, while steering it to believe an answer is incorrect (negative belief steering) strongly promotes retraction. That is to say, we can directly alter the model’s retraction behavior by intervening on this belief direction. Analysis of the steered models reveals two separate pathways through which internal beliefs control retraction. Negative belief steering first encourages the model to generate additional information (e.g., the entity’s birthplace) for verification rather than stopping immediately after the answer. Then, it also increases the model’s attention to answer tokens and refines their attention value vectors, which further promotes retraction.
@@
Finally, we show that the connection between the model’s belief and retraction holds for supervised fine-tuning (SFT). We fine-tune the model to retract wrong answers while maintaining commitment to correct ones. Consistent with prior work (Prakash et al., 2024; Ye et al., 2024; Muennighoff et al., 2025), we observe that straightforward SFT substantially improves in-distribution retraction performance: the model retracts more incorrect answers while still committing to correct ones. Beyond prior results, we demonstrate that the original belief direction continues to regulate retraction behavior. By probing the fine-tuned models, we show that SFT works by aligning the model’s internal belief more closely with factual correctness, leading to more accurate retraction decisions. This bridges mechanistic interpretability with training-based approaches, strengthening the robustness and generality of our findings.
@@
To summarize, our contributions are as follows: (1) We construct model-specific testbeds to evaluate an LLM’s retraction performance, and show that current LLMs can retract but do so only rarely. (2) We uncover a connection between a model’s internal belief and its external retraction behavior, and identify the underlying mechanism that governs this behavior. (3) We demonstrate that the causal influence of internal beliefs on retraction generalizes to supervised fine-tuned models, where more accurate beliefs lead to improved retraction performance.
@@ 2.1 Self-Correction in LLMs
A closely related concept to retraction is self-correction. Retraction can be viewed as an important step within self-correction but does not require producing a correct final answer as the goal. Previous work on self-correction primarily relies on multi-turn procedures, such as asking the model verification questions (Dhuliawala et al., 2024; Wu et al., 2024), prompting it to give feedback (Madaan et al., 2023; Zhang et al., 2024b; Liu et al., 2023), or directly instructing it to verify its initial responses (Kadavath et al., 2022; Yang et al., 2024c). By contrast, we study retraction as a spontaneous and immediate behavior, happening without explicit prompts to identify errors. This distinction is practically important, as users may not ask an LLM to re-check its answers. Relatively fewer studies have examined spontaneous self-correction (Ye et al., 2024; Zhao et al., 2025), showing that it can be acquired through elaborate training. Our work differs in that we explain how retraction emerges from the model’s internal representations, complementing training-based approaches.
2.2 Probing LLMs’ Beliefs
A series of studies leverages LLM’s internal representations to probe for truthfulness (Azaria and Mitchell, 2023; Marks and Tegmark, 2023b; Li et al., 2023; Liu et al., 2024). For example, Liu et al. (2024) propose the existence of a universal “truthfulness hyperplane” that separates true and false statements by training on a diverse collection of true-false datasets. However, many works (Li et al., 2023; Liu et al., 2024) evaluate probes only on synthetically constructed true-false claims, where they achieve high performance, but such settings may not reflect the distribution of hallucinations in real LLM outputs. Indeed, while some research demonstrates strong performance in detecting hallucinations on in-distribution, model-generated data (Azaria and Mitchell, 2023; CH-Wang et al., 2024; Orgad et al., 2024), these approaches often fail to generalize to out-of-distribution examples (Levinstein and Herrmann, 2023; Servedio et al., 2025).
In line with prior work, we train probes on external true-false datasets but evaluate them on the model’s own generated answers. Our findings provide a possible explanation for their limited generalization: such probes may not directly capture truth per se, but rather a model’s internal belief—its own judgments about the truth of the world (Levinstein and Herrmann, 2023; Schouten et al., 2024) during generation, which can diverge from ground-truth correctness. Crucially, we show that these belief signals are predictive of retraction behavior, suggesting that these probes may be tapping into dimensions of error awareness rather than factual truth itself.
@@ 3.1 Task Definition
Retraction denotes a model’s immediate acknowledgment that its generated answer is incorrect or does not fully satisfy the user’s requirements, regardless of whether it later produces a correct answer. To evaluate the retraction performance of current LLMs, we construct model-specific testbeds. We first collect questions from two knowledge-related datasets, Wikidata (e.g., “Name a writer who was born in Oran, Algeria”) and Celebrity (e.g., “Name a child of Joe Jackson”), which tend to elicit wrong answers, thereby creating a great opportunity to study retraction. Details of these two original datasets are provided in Appendix A.1.
Wikidata Celebrity
# Train # Test # Train # Test
Llama3.1-8B 1934 1202 1550 826
Qwen2.5-7B 1496 1072 1142
Olmo2-7B 1796 1260 1209
Table 1: Continuation dataset statistics. Note that Qwen2.5-7B and Olmo2-7B have no Celebrity training sets due to too few correct examples, which are used only for SFT in Section 6.
Continuation Dataset.
Based on the collected questions, we construct model-specific continuation datasets. Each example pairs a question with a model-generated answer, after which the model is prompted to continue generating to test whether it will retract, as illustrated below:
@@
To ensure that each incorrect answer is, in principle, correctable by the tested LLM, we first sample answers from the model via temperature decoding. For each answer, we create verification questions (e.g., “Where was {model’s answer} born?”; “What is {model’s answer}’s profession?”) and check whether the model’s responses to these questions conflict with the requirements of the original question, inspired by Dhuliawala et al. (2024). We retain two types of examples:
@@
  • Wrong Examples: The answer is factually incorrect, and the model’s responses to the verification questions contradict the original question, indicating that it should know the answer is incorrect.
We experiment with three popular LLMs from different model families, Llama3.1-8B-Instruct (Dubey et al., 2024, abbr. Llama3.1-8B), Qwen2.5-7B-Instruct (Yang et al., 2024a, abbr. Qwen2.5-7B), and Olmo2-1124-7B-Instruct (OLMo et al., 2025, abbr. Olmo2-7B). The data statistics are listed in Table 1. See Appendix A.2 for details. In the following sections, we use Wikidata and Celebrity to denote the model-specific continuation datasets instead of the original datasets.
@@
We use Llama3.3-70B-Instruct as a judge (Zheng et al., 2023) to automatically assess whether the tested model retracts the given answer in its response. See Appendix A.6 for details and reliability analyses. We then calculate the following two metrics to evaluate the model’s retraction performance:
@@
[…]
$|\text{Wrong}|$ denotes the number of wrong examples, and $|\text{Retraction}|$ indicates the number of examples that the tested model retracts according to the judgment of Llama3.3-70B-Instruct. Higher retraction recall and precision together represent better retraction performance.
@@ 3.2 Models Can Retract, but Do So Infrequently
Wikidata Celebrity
Prec. Rec. Prec. Rec.
Llama3.1-8B 0.9012 0.2579 0.7722 0.1477
Qwen2.5-7B 0.8824 0.1119 0.9667 0.0290
Olmo2-7B 0.9881 0.1317 0.8824 0.0150
@@
As shown in Table 2, models consistently exhibit low but non-zero retraction recall on our datasets. We infer that LLMs have the capability to retract incorrect answers, but the consistently low recall (at most $25\%$) highlights that such retractions are rare. Recall that our verification questions provide clear evidence that the model knows that the incorrect answers in our datasets are indeed incorrect. Thus, the model appears to have both the knowledge and the ability to retract. Then, why do LLMs fail to retract more incorrect answers? What factors govern their retraction behavior?
@@ 4 Model Belief Guides Retraction
4.1 Probing for Belief
To investigate the gap between LLMs’ parametric knowledge measured by factoid questions and their failure to retract incorrect answers, we build on prior work that probes internal representations of truthfulness (Azaria and Mitchell, 2023; Marks and Tegmark, 2023b; Li et al., 2023). Here, we use the term internal belief rather than truthfulness to emphasize the distinction between a model’s internal assessment of correctness and ground-truth correctness. Our key question is: when the model’s parametric knowledge implies that its answer is wrong, does its internal representation reflect this during answer generation?
@@ 4.1 Probing for Belief
To prevent overfitting to a single dataset, we follow Liu et al. (2024) and train our probes on a diverse set of external true-false datasets, including 800 examples each from Natural Questions (Kwiatkowski et al., 2019), Trivia QA (Joshi et al., 2017), and SciQ (Welbl et al., 2017). All three are short-answer, closed-book QA tasks with a format similar to Wikidata and Celebrity. Each dataset is balanced with a 50/50 split of correct and incorrect answers, where the incorrect answers are generated by GPT-4-turbo. We denote this collection as Universal Truthfulness QA (UTQA) dataset.
@@
Figure 2: Layer-wise AUROC of belief scores for predicting factual correctness and retraction in Llama3.1-8B. An AUROC of 0.5 corresponds to random guessing. Results are averaged over three runs with different random seeds, and error bars denote standard deviation.
Probe Setup.
For each LLM layer, we train a linear probe on the UTQA dataset using the hidden states after the given answer. These probes learn to distinguish correct and incorrect answers on UTQA and thus serve as proxies for the model’s internal belief. We then apply the probes to Wikidata and Celebrity examples to investigate how the model’s internal belief relates to both factual correctness and retraction behavior. To quantify the relationship, we report AUROC (Area Under the Receiver Operating Characteristic Curve), treating belief scores as the decision score and either binary factual correctness or retraction labels as the ground truth. Because internal belief and retraction are hypothesized to be negatively correlated, we use 1belief score1-$\text${belief score} when predicting retraction. A higher AUROC indicates that belief scores are reliable predictors of the target label, while an AUROC of 0.5 implies no discriminative power.
Results.
From Figure 2, we can see that belief probes are less predictive of factual correctness but much more reliable for predicting retraction.
(1) Since factual correctness on our test sets reflects the model’s parametric knowledge as assessed by verification questions, the suboptimal AUROC of belief scores indicates a misalignment between the model’s momentary internal belief and its stored knowledge. This further verifies LLMs’ limitations in knowledge manipulation (Allen-Zhu and Li, 2025; Berglund et al., 2024), from the perspective of internal representations.
(2) More importantly, we find that the model’s internal belief, although obtained without retraction-related data, is a better indicator of whether the model retracts its own generated answers, with predictive power emerging in the middle layers. In particular, low belief scores correspond to retraction. [^2] Internal belief is conceptually distinct from uncertainty. Appendix B.2 shows that uncertainty is a much weaker predictor of retraction. This suggests that the probed direction captures the model’s internal judgment at that moment instead of objective truth, and is manifested in its subsequent behavior (i.e., retracting or committing).
These results generalize across model families, including Qwen2.5-7B and Olmo2-7B, and extend to larger models such as Llama3.1-70B-Instruct, as shown in Appendix B.1.
4.2 Steering Internal Belief Affects Retraction
Our probing results establish a correlation between the model’s internal belief and its retraction behavior. To demonstrate that internal belief causally influences retraction behavior, we steer the model’s hidden states towards positive belief (i.e., believe an answer is correct) and negative belief (i.e., believe an answer is incorrect) directions.